OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 1 over 4(g − 16)

OpenStudy (zzr0ck3r):

a(b-c) = ab-ac

OpenStudy (zzr0ck3r):

now try

OpenStudy (anonymous):

is it a expression or term

OpenStudy (anonymous):

its an expression

OpenStudy (zzr0ck3r):

2(x+4) = 2(x)+2(4) = 2x+8 now you try

OpenStudy (zzr0ck3r):

2(x-4) = 2(x) - 2(4) = 2x-8 now you try

OpenStudy (anonymous):

can you write in expression form

OpenStudy (anonymous):

I really dont know D: